# Conforming Spline-Based Parameterisation Techniques for Plane Graphs

### v1

@@ 3 Face Parameterisation
wherein $p\geq 2$ denotes the polynomial order while $N_{0}$ tunes the initial spacing of the knot spans. In practice, we utilise $N_{0}=5$. We note that, albeit uncommon, interior knots may also be repeated a maximum of $p-1$ times in (36) without violating the assumption that each spline fit $s_{q}\in C^{1}([0,1],\mathbb{R}^{2})$.
@@
$\min\limits_{s\in\mathcal{V}_{\xi}^{2}(\Xi_{i})}\quad\frac{1}{N_{i}}\sum_{j}\left\|s(\xi_{j})-p_{j}\right\|^{2}+\lambda\int\limits_{(0,1)}\left\|\frac{\partial^{2}s}{\partial\xi^{2}}\right\|^{2}\mathrm{d}\xi,$
@@
Given the quadrangular faces $q\in\mathcal{Q}_{i}$, the unique knots assigned to the $F_{i}\ni e\rightarrow\hat{e}\in\partial E_{i}$ are propagated into the interior as in Figure 9(a). Depending on the layout, some interior edges’ knotvectors may not follow from the boundary edges. In this case, we assign a uniform knotvector $\Xi^{I}$ (with the property that $\Xi^{I}=-\Xi^{I}$) to the isolated edges. The number of interior knots is chosen based on the average number of knots in all other knotvectors. For an example, see Figure 9(b).
Figure 9: Figure showing how the knotvectors $\Xi=w^{\Xi}_{i}(\hat{e}),\,\hat{e}\in\partial E_{i}$ are propagated from the boundary edges into the interior to form the enriched weight function $w^{\Xi}_{i}(\,\cdot\,)$ that assigns the same knotvector to neighboring edges of the $q\in\mathcal{Q}_{i}$ by taking the union of all boundary knotvectors that align with a pair of edges in the quad layout (a). The result is a knotvector pair $(\Xi_{\mu}^{j},\Xi_{\nu}^{j})$ for each quad face $q_{j}\in\mathcal{Q}_{i}$ / patch ${\hat{\Omega}}_{j}^{\square}\subset{\hat{\Omega}}_{i}$ which conforms with neighboring cell’s knotvectors. This results in the coarsest element segmentation that enables the creation of a basis $\mathcal{V}_{h,i}\subset H^{1}({\hat{\Omega}}_{i})$ that is compatible with the boundary correspondence $\mathbf{f}_{i}:\partial{\hat{\Omega}}_{i}\rightarrow\partial{\Omega}^{s}_{i}$ in the sense that the set $\mathcal{U}^{\mathbf{f}_{i}}\neq\emptyset$. Figure (b) shows an example of a template $\mathcal{T}_{i}$ where not all interior knotvectors follow from the boundary knotvectors (red), in which case a suitable knotvector is selected automatically.
The figure also illustrates how the use of dyadically refined knotvectors avoids the introduction of a dense element subdivision when taking the union of two or more knotvectors. Arbitrarily-placed knots may furthermore lead to minuscule knotspans upon taking the union of several knotvectors, which, in turn, leads to conditioning issues of matrices assembled over the associated spline space. While not immediately evident, we emphasise that potential conditioning issues make the use of dyadically refined knotvectors essentially mandatory in practice.
@@
where $\mathbf{m}^{j}:{\Omega}^{\square}\rightarrow{\hat{\Omega}}_{j}^{\square}$ is the bilinear diffeomorphic map between ${\Omega}^{\square}$ and the $j$-th patch ${\hat{\Omega}}^{\square}_{j}\subseteq{\hat{\Omega}}$. We note that the canonical basis $\{\mathcal{V}_{h}^{\text{disc}}\}$ of $\mathcal{V}_{h}^{\text{disc}}\subset L^{2}({\hat{\Omega}}_{i})$ forms a partition of unity on ${\hat{\Omega}}$. We may define the subspace $\mathcal{V}_{h}:=\mathcal{V}_{h}^{\text{disc}}\cap C^{0}({\hat{\Omega}})$, with $\mathcal{V}_{h}\subset H^{1}({\hat{\Omega}})$, and the set $\mathcal{U}^{\mathbf{F}}_{h}:=\{\boldsymbol{\phi}_{h}\in\mathcal{V}_{h}^{2}\,\,|\,\,\boldsymbol{\phi}_{h}=\mathbf{F}\text{ on }\partial{\hat{\Omega}}\}$. By construction, the set $\mathcal{U}^{\mathbf{f}}$, that takes $\mathbf{F}$ as the spline-based boundary correspondence $\mathbf{f}:\partial{\hat{\Omega}}\rightarrow\partial{\Omega}^{S}$, satisfies $\mathcal{U}^{\mathbf{f}}\neq\emptyset$.
@@
are given by the union of the images of the $s=w^{S}(e),\,e\in E$. As such, they are fully conforming, both in the sense of watertightness as well as parameterically. We note, however, that if two faces $F_{i}$ and $F_{j}$ have shared edges, the $s=w^{S}(e)$, with ±$\pm e\in F_{i}$ and $\mp e\in F_{k}$, may be expressed in differing supersets of the canonical knotvector $\Xi=w^{\Xi}(e)$ of $s$ under the two maps $\mathbf{x}_{h,i}$ and $\mathbf{x}_{h,j}$ (see Figure 10(a)). More precisely, let $e\in F_{i}$ with $-e\in F_{j}$ be a shared edge. Furthermore, let $\hat{e}_{1}$ and $-\hat{e}_{2}$ be the associated local edges under the correspondences $F_{i}\rightarrow\partial E_{i}$ and $F_{j}\rightarrow\partial E_{j}$, respectively. Then, generally, $w^{\Xi}_{j}(\hat{e}_{1})\neq w^{\Xi}_{k}(-\hat{e}_{2})$. If desired, this can be computationally inexpensively remedied by prolonging neighboring faces to a unified element segmentation, as in Figure 10(b).
@@ 4 Results
Applying the heuristic strategy 1, we obtain the TSG depicted in Figure 11 (c) and the finalised spline-based parameterisation of the input graph in Figure 11 (a). Figure 11 (b) shows the finalised parameterisation after selection based on strategy 2. Both selection strategies yield similar results with the exception of the face marked by the letter A, for which strategy 2 selects a template that results in a parameterisation with notably improved uniformity. However, both spline parameterisations’ faces are all bijective without the need for untangling.
@@
Figure 14: Spline-based parameterisation of a plane graph representing the African continent with template selection based on (29) with template optimisation based on (2.2).
@@ Acknowledgements
Furthermore, the authors are grateful for the coding help they received from the Nutils core development team.

### v2

@@ 3 Face Parameterisation
wherein $p\geq 2$ denotes the polynomial order while $N_{0}$ tunes the initial spacing of the knot spans. We note that, albeit uncommon, interior knots may also be repeated a maximum of $p-1$ times in (34) without violating the assumption that each spline fit $s_{q}\in C^{1}([0,1],\mathbb{R}^{2})$.
@@
$\min\limits_{s\in\mathcal{V}_{\xi}^{2}(\Xi^{i})}\quad\frac{1}{N_{i}}\sum_{j}\left\|s(\xi_{j})-p_{j}\right\|^{2}+\lambda\int\limits_{(0,1)}\left\|\frac{\partial^{2}s}{\partial\xi^{2}}\right\|^{2}\mathrm{d}\xi,$
@@
Given the quadrangular faces $q\in\mathcal{Q}_{i}$, the unique knots assigned to the $F_{i}\ni e\rightarrow\hat{e}\in\partial E_{i}$ are propagated into the interior as in Figure 14. Depending on the layout, some interior edges’ knotvectors may not follow from the boundary edges. In this case, we assign a uniform knotvector $\Xi^{I}$ (with the property that $\Xi^{I}=-\Xi^{I}$) to the isolated edges. The number of interior knots is chosen based on the average number of knots in all other knotvectors. For an example, see Figure 14, right.
Figure 14: Figure showing how the knotvectors $\Xi=w^{\Xi}(e)$ (blue) are propagated from the global edge $e\in F_{i}$ to the associated local edge $\hat{e}=\phi_{i}^{-1}(e)$ while simultaneously taking the union of opposite knots to create a local weight function $w_{i}^{\Xi}(\,\cdot\,)$ (red). The result is a knotvector pair $(\Xi_{\mu}^{j},\Xi_{\nu}^{j})$ for each quad face $q_{j}\in\mathcal{Q}_{i}$ / patch ${\hat{\Omega}}_{j}^{\square}\subset{\hat{\Omega}}_{i}$ which conforms with neighboring cell’s knotvectors (center). This results in the coarsest element segmentation that enables the creation of a basis $\mathcal{V}_{h,i}\subset H^{1}({\hat{\Omega}}_{i})$ that is compatible with the boundary correspondence $\mathbf{f}_{i}:\partial{\hat{\Omega}}_{i}\rightarrow\partial{\Omega}^{s}_{i}$ in the sense that the set $\mathcal{U}^{\mathbf{f}_{i}}\neq\emptyset$. The right figure shows an example of a template $T_{i}$ where not all interior knotvectors follow from the boundary knotvectors, in which case a suitable knotvector is selected automatically.
The figure additionally illustrates how the use of dyadically refined knotvectors avoids the introduction of a dense element subdivision when taking the union of two or more knotvectors. Arbitrarily-placed knots may furthermore lead to minuscule knotspans upon taking the union of several knotvectors, which, in turn, leads to conditioning issues of matrices assembled over the associated spline space. While not immediately evident, we emphasise that potential conditioning issues make the use of dyadically refined knotvectors essentially mandatory in practice.
@@
where $\mathbf{m}^{j}:{\Omega}^{\square}\rightarrow{\hat{\Omega}}_{j}^{\square}$ is the bilinear diffeomorphic map between ${\Omega}^{\square}=(0,1)^{2}$ and the $j$-th patch ${\hat{\Omega}}^{\square}_{j}\subseteq{\hat{\Omega}}$. We note that the canonical basis $\{\mathcal{V}_{h}^{\text{disc}}\}$ of $\mathcal{V}_{h}^{\text{disc}}\subset L^{2}({\hat{\Omega}}_{i})$ forms a partition of unity on ${\hat{\Omega}}$. We may define the subspace $\mathcal{V}_{h}:=\mathcal{V}_{h}^{\text{disc}}\cap C^{0}({\hat{\Omega}})$, with $\mathcal{V}_{h}\subset H^{1}({\hat{\Omega}})$, and the set $\mathcal{U}^{\mathbf{F}}_{h}:=\{\boldsymbol{\phi}_{h}\in\mathcal{V}_{h}^{2}\,\,|\,\,\boldsymbol{\phi}_{h}=\mathbf{F}\text{ on }\partial{\hat{\Omega}}\}$. By construction, the set $\mathcal{U}^{\mathbf{f}}$, that takes $\mathbf{F}$ as the spline-based boundary correspondence $\mathbf{f}:\partial{\hat{\Omega}}\rightarrow\partial{\Omega}^{S}$, satisfies $\mathcal{U}^{\mathbf{f}}\neq\emptyset$.
@@
are given by the union of the images of the $s=w^{S}(e),\,e\in E$. As such, they are fully conforming, both in the sense of watertightness as well as parameterically. We note, however, that if two faces $F_{i}$ and $F_{j}$ have shared edges, the $s=w^{S}(e)$, with ±$\pm e\in F_{i}$ and $\mp e\in F_{k}$, may be expressed in differing supersets of the canonical knotvector $\Xi=w^{\Xi}(e)$ of $s$ under the two maps $\mathbf{x}_{h,i}$ and $\mathbf{x}_{h,j}$ (see Figure 15(a)). More precisely, let $e\in F_{i}$ with $-e\in F_{j}$ be a shared edge. Furthermore, let $\hat{e}_{1}$ and $-\hat{e}_{2}$ be the associated local edges under the correspondences $\phi_{i}^{-1}:F_{i}\rightarrow\partial E_{i}$ and $\phi_{j}^{-1}:F_{j}\rightarrow\partial E_{j}$, respectively. Then, generally, $w^{\Xi}_{j}(\hat{e}_{1})\neq w^{\Xi}_{k}(-\hat{e}_{2})$. If desired, this can be computationally inexpensively remedied by prolonging neighboring faces to a unified element segmentation, as in Figure 15(b).
@@ 4 Results
Applying the heuristic strategy 1, we obtain the TSG depicted in Figure 12 (c) and the finalised spline-based parameterisation of the input graph in Figure 16 (a). Figure 16 (b) shows the finalised parameterisation after selection based on strategy 2. Both strategies yield similar results with the exception of the face marked by the letter A, for which strategy 2 selects a template that results in a parameterisation with notably improved uniformity. However, both spline parameterisations’ faces are all bijective without the need for untangling.
@@
Figure 19: Spline-based parameterisation of a plane graph representing the African continent with template selection based on (27) and template optimisation based on (2.2).
@@ 6 Conclusion
The authors have no relevant financial or non-financial interests to disclose. The authors have no competing interests to declare that are relevant to the content of this article. All authors certify that they have no affiliations with or involvement in any organization or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript. The authors have no financial or proprietary interests in any material discussed in this article.
Appendix A Floater’s Algorithm
Floater’s algorithm (see Section 1.5) is a computationally inexpensive method to acquire a decent approximation of the harmonic map $\mathbf{x}^{-1}:\Omega^{S}\rightarrow{\hat{\Omega}}^{\mathbf{r}}$ over a polygonal domain $\Omega^{S}_{h}\approx\Omega^{S}$. The approach samples ordered point sets $X=\left(x_{1},\ldots,x_{n}\right)$ and $\hat{X}=\left(\hat{x}_{1}^{r},\ldots,\hat{x}_{i}^{r}\right)$ from $\partial\Omega^{S}$ and $\partial{\hat{\Omega}}^{\mathbf{r}}$, respectively, leading to a canonical discrete correspondence $(\mathbf{f}_{h}^{\mathbf{r}\rightarrow\mathbf{x}})^{-1}:X\rightarrow\hat{X}$, with $(\mathbf{f}_{h}^{\mathbf{r}\rightarrow\mathbf{x}})^{-1}(x_{i})=\hat{x}_{i}^{r}$. The boundary of the polygon $\Omega_{h}^{S}$ is forwarded to a meshing routine which returns the triangulation $\mathcal{T}_{h}$ with boundary vertices $X\subset\mathbb{R}^{2}$ and a total number of $N>n$ vertices. An approximately harmonic map is then reconstructed from the canonical piecewise-linear $C^{0}$-continuous Lagrange basis $P_{h}$ over $\mathcal{T}_{h}$, which leads to a linear problem of the form $Ac=Bf^{r}$. Here, the solution vector $c\in\mathbb{R}^{2(N-n)}$ contains the approximation’s weights (w.r.t. to $P_{h}\times P_{h}$) that are not fixed from the boundary correspondence, while the vector $f^{r}\in\mathbb{R}^{2n}$ is the concatination of the $\hat{x}_{i}^{r}\in\hat{X}$. For a computationally inexpensive method to assemble the SPD $A\in\mathbb{R}^{2N\times 2N}$ and the sparse $B\in\mathbb{R}^{2N\times 2n}$, we refer to [9]. The approximation $\left(\mathbf{x}_{h}^{\mathbf{r}}\right)^{-1}\approx\left(\mathbf{x}^{\mathbf{r}}\right)^{-1}$ maps the triangulation $\mathcal{T}_{h}$ onto a triangulation $\hat{\mathcal{T}}_{h}^{r}$ of ${\hat{\Omega}}^{\mathbf{r}}_{h}\approx{\hat{\Omega}}^{\mathbf{r}}$ which automatically yields a piecewise linear approximation $\mathbf{x}_{h}^{\mathbf{r}}:{\hat{\Omega}}^{\mathbf{r}}_{h}\rightarrow{\Omega}_{h}^{S}$ of an inversely harmonic map from the correspondence of the vertices $\hat{\mathcal{T}}^{r}_{h}\ni\hat{v}_{i}\rightarrow v_{i}\in\mathcal{T}_{h}$.
Appendix B Template Creation using Patch Adjacency Graphs
over all interior vertices $V^{\mathbf{r}}\ni v^{\mathbf{r}}\in V^{\mathbf{r}}_{\text{int}}$. Here, $\operatorname{ReLU}(\,\cdot\,)$ denotes the ReLU activation function $\operatorname{ReLU}(x):=\max\{0,x\}$, while $\tau\geq 0$ is an offset. We note that $C(\,\cdot\,,\tau)$ vanishes once all $\boldsymbol{\nu}(v^{\mathbf{r}},q^{\mathbf{r}})$ are above the threshold $\tau$ and the minimisation is terminated. Minimisation based on (52) is nonconvex and the cost function is nonsmooth. However, a gradient-based algorithm nevertheless converges after a small number of iterations as long as the initial guess represents a layout for which all $\boldsymbol{\nu}(v^{\mathbf{r}},q^{\mathbf{r}})$ are close to or exceed $\tau$. Starting on $\tau=0$, its value is incremented in an outer loop until optimisation no longer yields a minimiser for which $C(\mathcal{T}^{\mathbf{r}},\tau)=0$. In practice, the optimisation terminates in a fraction of a second thanks to the low number of interior vertices. Reusing the previously computed map ${\mathbf{r}}_{h}:{\hat{\Omega}}\rightarrow{\hat{\Omega}}^{\mathbf{r}}_{h}$ to create an initial layout for optimisation based on (52), in this way, hundreds of templates can be optimised in a matter of seconds.
In practice, the final step is only carried out for a single template which is selected from a large catalogue of optimised templates.
All required steps are summarised by Figure 25.